Jenny can row 20 miles with the current in 2 hours, or 12 miles against the current in 3 hours. Find the rate of Jenny's boat in still water and the rate of the current.
Answer by mananth(16946)

rowing speed =x mph
current speed =y mph
against current 3 hours
with current 2 hours

Distance with current 20 miles distance against current 12
t=d/r against current (x-y)
12.00 / ( x - y )= 3.00

3.00 x - -3.00 y = 12.00 ....................1
with current (x+y)
20.00 / ( x + y )= 2.00
2.00 ( x + y ) = 20.00
2.00 x + 2.00 y = 20.00 ...............2
Multiply (1) by 2.00
Multiply (2) by 3.00
we get 2.00
6.00 x + -6.00 y = 24.00
6.00 x + 6.00 y = 60.00
12.00 x = 84.00
/ 12.00
x = 7 mph

plug value of x in (1) y
3 x -3 y = 12
21 -3 -21 = 12
-3 y = 12
-3 y = -9 mph
y = 3
rowing 7 mph
current 3 mph